Find a fraction in which the difference of the numerator and the denominator is 2 and sum is 8.
Answer by Theo(13342)

let x be the numerator.
then x+2 = the denominator.

x + x + 2 = 8 becomes:
2x + 2 = 8 becomes:
2x = 6 becomes:
x = 3

your fraction is 3/5.